# Demystifying Domain-adaptive Post-training for Financial LLMs

### v1

@@ 2.1 FinDaP Capabilities
$\bullet$ Others. Additionally, domains may vary significantly in their sensitivity. For instance, the medical domain is highly sensitive, requiring utmost accuracy and strict adherence to ethical considerations. In contrast, domains such as entertainment may have more relaxed requirements. Another important consideration is multi-modality, as some domains require handling multiple types of input and output formats. For example, the healthcare domain may involve processing medical images alongside textual reports, while the e-commerce domain may integrate product descriptions, images, and customer reviews into a unified response. Similarly, scientific research often combines charts, graphs, and textual analysis to present findings effectively. While we acknowledge these additional aspects, we leave those for future work and concentrate on the four primary capabilities discussed above.
2.2 FinDaP Evaluation
With the above breakdown of capabilities, our evaluation framework consists of a suite for assessing these capabilities using development sets and unseen (held-out) evaluation sets. Our development set is directly split from the training data at each stage. Table 1 outlines the capabilities and the evaluation benchmarks selected to cover these capabilities. Crucially, we did not examine scores on our unseen set while developing the models, which allows us to observe how much we may have overfitted to particular evaluations in our decisions around training recipe. For the unseen tasks (Table 1), we manually review each individual dataset and have the following considerations.
@@ 2.2 FinDaP Evaluation
ESG Issue Classification MLESG 300 Chen et al. (2023b)
@@
Table 1: Summary of our evaluation dataset. New datasets released with FinDaP are color-highlighted for emphasis.
@@
$\bullet$ Benchmarking tasks. Corresponding to the capabilities, we consider a diverse set of benchmarking tasks. For concepts, we include knowledge tasks in the general domain, such as AI2-ARC (Clark et al., 2018), as well as in finance, such as MMLU-Finance (Hendrycks et al., 2021). For tasks, we consider general tasks, such as Social-IQA (Welbl et al., 2017), and domain-specific tasks, such as MLESG (Chen et al., 2023b). Notably, we intentionally include a few financial tasks such as Flare-TATQA (Zhu et al., 2021) and SM-Bigdata (Soun et al., 2022) that require understanding of tabular data, as this data format is common in this domain. For IF/Chat capabilities, we utilize popular instruction-following benchmarks, such as MT-Bench (Zheng et al., 2023). For reasoning, we include general reasoning tasks, such as MathQA (Amini et al., 2019a) and Hellaswag common sense reasoning (Zellers et al., 2019), as well as domain-specific reasoning tasks, such as CRA-ProroSeguro claim analysis (Feng et al., 2024). We also construct a new benchmark on CFA-Challenge based on CFA Level III, one of the most challenging financial exams that requires comprehensive reasoning (Khamnuansin et al., 2024; Callanan et al., 2024).
@@
$\bullet$ Evaluation method. We split our evaluation set into two types based on their exposure to Instruction tuning (IT) data (Table 1). The first type, Similar, includes tasks whose types have been encountered during training, even if the specific tasks themselves are unseen (e.g., a new NER task). The second type, Novel, includes tasks whose types have not been seen during training, representing entirely new challenges for the model (e.g., stock movement prediction). We use two different evaluation methods based on the nature of the benchmarks. For knowledge and NLP tasks (e.g., NER), we employ a straightforward direct answer evaluation. For reasoning tasks (e.g., CFA-Challenge), we use a 0-shot chain-of-thought (CoT) (Wei et al., 2023) answer evaluation to enhance the reliability of our evaluation. This also exposes the reasoning path, allowing us to investigate the causes of incorrect answers and enabling a more fine-grained comparison across different models.
2.3 Comparison with Existing Finance LLMs
Table 2 compares Llama-Fin with popular open finance-specific LLMs: FinTral (Bhatia et al., 2024), PIXIU (Xie et al., 2023), FinLLM (Xie et al., 2024b), and AdaptLLM (Cheng et al., 2024). Among these models, only FinTral incorporates all three training stages: CPT, IT, and PA. However, unlike Llama-Fin, FinTral and other models lack a systematic design and fail to include carefully crafted components, as illustrated in Figure 1. For instance, none of these models have an in-depth analysis of design choices for different post-training stages, let alone introduce a novel approach for constructing preference data.
Finance LLM Capability Training Evaluation
Continue Pre-training Instruction Tuning Preference Alignment Task Method
General Domain Target Domain General Task Domain Task On-policy General Task Domain Task Reasoning Task Direct Answer CoT Answer
ORM PRM
FinTral Concept, Task
PIXIU Task
FinLLM Concept, Task
AdaptLLM Concept
Llama-Fin
Concept, IF/Chat,
Task, Reasoning
Table 2: Comparison between Llama-Fin with other open finance LLMs.
3 Model Training
As shown in Figure 1, FinDaP’s training starts with curated text and prompts (§3.1), followed by joint Continual Pre-training and Instruction Tuning (§3.4), and then Preference Alignment (§3.5).
3.1 Data
Text Curation. To introduce domain concepts while preserving general concepts, we curate texts for CPT. Table 3 summarizes the texts curation datasets. Specially, for general concepts, research has shown that a ‘small’ amount of general text (as little as 1%) can effectively mitigate the forgetting issue (Scialom et al., 2022). Therefore, we focus on collecting a relatively small but high-quality set of general-domain text. To achieve this, we use verifiable text, which is text written by humans and previously used in supervised tasks in the literature. Note that this contrasts with using unverifiable web text such as C4 (Raffel et al., 2020).
Capability Domain CPT Dataset Size Reference
Concept General NaturalInstrution 100,000 Mishra et al. (2022)
PromptSource 100,000 Bach et al. (2022)
Math 29,837 Amini et al. (2019b)
Aqua 97,500 Ling et al. (2017)
CREAK 10,200 Onoe et al. (2021)
ESNLI 549,367 Camburu et al. (2018)
QASC 8,130 Khot et al. (2020)
SODA 1,190,000 Kim et al. (2022)
StrategyQA 2,290 Geva et al. (2021)
UnifiedSKG 779,000 Xie et al. (2022)
GSM8K 7,470 Cobbe et al. (2021)
ApexInstr 1,470,000 Huang et al. (2024b)
DeepmindMath 379,000 Saxton (2019)
DialogueStudio 1,070,000 Zhang et al. (2023)
Finance Fineweb-Fin 4,380,000 -
Book-Fin 4,500 -
Total 10,177,294
Table 3: Summary of curated texts. New datasets released with FinDaP are color-highlighted for emphasis.
For domain concept, our goal is to collect both a large volume of data and maintain high quality. Following practices from the literature on training general LLMs (Lambert et al., 2024; Gunasekar et al., 2023), we source financial texts from primarily relevant websites and books. Details on the collection process can be found in Appendix A.
Capability Domain Task IT Dataset Size Reference
Tasks Finance Relation Cls. FingptFinred 27,600 Sharma et al. (2022)
NER FingptNERCls 13,500 Yang et al. (2023)
FingptNER 511 Alvarado et al. (2015)
Headline Cls. FingptHeadline 82,200 Sinha et al. (2020)
Sentiment Cls. SentimentCls 47,600 Yang et al. (2023)
SentimentTra 76,800 Yang et al. (2023)
Summariz. TradeTheEvent 258,000 Zhou et al. (2021)
IF/Chat General IF/Chat SelfInstruct 82,000 Wang et al. (2022)
SlimOrca 518,000 Lian et al. (2023)
UltraChat 774,000 Ding et al. (2023)
ShareGPT 100,000 Link
Finance QA FinanceInstruct 178,000 Link
FingptConvfinqa 8,890 Chen et al. (2022)
FlareFinqa 6,250 Chen et al. (2021)
FlareFiqa 17,100 Yang et al. (2023)
Reasoning Math QA OrcaMath 200,000 Mitra et al. (2024)
MetaMathQA 395000 Yu et al. (2023)
MathInstruct 262,000 Xiang Yue (2023)
Code QA MagicodeInstruct 111,000 Luo et al. (2023)
Finance CFA Exam Exercise 2,950 -
Total 3,161,401
Table 4: Summary of our curated prompts. New datasets released with FinDaP are color-highlighted for emphasis. For datasets without formal references but only a URL, we provide their links.
Prompt Curation. Prompts represent the diverse ways users may interact with models and serves the essential component for IT and PA. Table 4 summarizes the prompts curation datasets. Specifically, we conduct a broad survey and source general, financial, instruction-following, and reasoning tasks from public datasets. To promote diversity, we include datasets like Flare-FinQA  (Chen et al., 2021), a large open QA dataset in finance, and UltraChat (Ding et al., 2023), a dataset shown to perform well for IT in the literature (Tunstall et al., 2024; Ivison et al., 2024). Additionally, we find that exercises or demonstrations from books that were curated in §3.1 is valuable for reasoning tasks as they usually involve challenging reasonings and come with ground truth answers and sometimes even include human-written chain-of-thought (CoT) explanations. Details on the prompts used to extract these exercises can be found in Appendix B.
3.2 Continual Pre-training (CPT)
Figure 2: Average performance on selected datasets for training Llama3-8b-instruct on our CPT-In, CPT-Gen and CPT-Mix. The selected datasets are chosen for illustration purpose based on their ability to illustrate the general trend.
From Text to CPT Data. A key challenge in CPT is what kind of data we should use. Given the general and domain-specific texts introduced in §3.1, we can construct three versions of CPT data, CPT-In contains only the financial (in-domain) text, CPT-Gen contains only the general domain data, and CPT-Mix contains the mixture of the CPT-In and CPT-Gen.
Key Data Experiments. We conduct CPT on each of the three versions of data. As shown in Figure 2, we observe that while CPT-In and CPT-Gen outperforms in financial (Fig 2a) and general (Fig 2b) tasks, respectively, CPT-Mix achieves the best overall. This is expected as CPT-In can cause catastrophic forgetting on the general tasks, while incorporating general domain concepts in CPT-Mix acts as ‘replay’ mechanism to mitigate it (Scialom et al., 2022). We can also see that none of the CPT-trained LLMs outperform their base. This is unexpected because CPT invovles post-training on more specialized data, which should enhance the performance. By analyzing the output, we attribute this issue to the model forgetting how to follow instructions effectively after CPT. To quantify this finding, we evaluate the instruction following ability of these models using MT-Bench. The two-turn average scores for CPT-Mix, CPT-In, and CPT-Gen are 1, 1, and 1.0125, respectively, while the base model, achieves a score of 7.8875. These confirm that the conventional CPT applied to a instruction-tuned LLM can cause serious forgetting on instruction-following (IF) capability. In §3.3, we will see how jointly train IT and CPT can help mitigate such forgetting issue.
3.3 Instruction Tuning (IT)
From Prompt to IT Data.
Figure 3: Average performance on selected datasets for training Llama3-8b-instruct on our IT-In, IT-Gen and IT-Mix.
We introduced our prompt curation in §4. We create the responses for IT by filtering existing responses or creating new responses. For prompts with existing responses, we generally keep the original responses if they were written by a human or a strong model, such as GPT-4. We also filter out empty responses. For prompts without responses, for example, exercises extracted from books that may not have solutions provided, we generate new responses using GPT-4o. Similar to CPT data, we construct three versions of IT data, IT-In, which contains only financial (in-domain) tasks, IT-Gen, which contains only general tasks, and IT-Mix, which includes a mixture of the IT-In and IT-Gen.
Key Data Experiments. Similar to CPT, we conduct IT to each of the three versions. From Figure 3, we observe that unlike CPT, forgetting is significantly reduced. Specifically, all versions of IT are no longer worse than their base versions, indicating that the ability to follow instructions is not as severely forgotten as in CPT. This is further supported by the MT-Bench scores, where we obtained 7.2031, 6.2094, and 7.3219 for IT-Mix, IT-In and IT-Gen, repsetively, all of which are significantly better than the CPT counterparts.
Figure 4: Average performance on selected datasets for training Llama3-8b-instruct on IT-Mix with full-model finetuning (IT-Mix) and LoRA finetuning (IT-Mix (LoRA)).
Comparison with LoRA. Another popular approach to adapt the LLM to specific domain is Parameter-efficient Fine-tuning (PEFT), where the LLM parameters remain fixed, and only a small set of additional parameters are trained. This approach naturally mitigates forgetting issues and is more efficient in terms of trainable parameters. However, whether it can achieve performance comparable to full-model training is unclear. In Figure 4, we experiment with PEFT, specifically using LoRA (Hu et al., 2021), with a rank size of 128, [^2] Further decreasing or increasing the rank size did not show improvement in our preliminary experiments. For example, rank size of 32, 128 and 512 yield overall averages across 10 general tasks of 0.5267, 0.5331, and 0.5215, respectively, showing only minor differences. and compare its performance with full-model finetuning (IT-Mix). We observe that with and without LoRA performs similarly, confirming that LoRA is effective for task adaptation. However, the novel tasks still show little improvement, highlighting that task generalization still remains a significant challenge.
Figure 5: Average performance on selected datasets for training Llama3-8b-instruct on CPT-Mix and IT-Mix jointly (CPT-Mix + IT-Mix) and sequentially (CPT-Mix $\rightarrow$ IT-Mix).
3.4 Combining CPT and IT
A natural choice is to conduct CPT and IT sequentially (Lambert et al., 2024). On the one hand, this is flexible as it allows for different settings (e.g., data size) in each stage. On the other hand, it does not help prevent forgetting during the CPT stage, leaving the LLM dependent on IT to ‘recover’ its instruction-following capability. To make a more grounded decision, we conduct experiments on both sequential and joint training approaches. In joint training, an additional hyperparameter to consider is the mixture ratio. We find that down-sampling CPT data to match the size of IT data is the most effective strategy. Results for other strategies, such as no-sampling are provided in Appendix D.
Figure 6: Average performance on selected datasets for PEFT or full model fine-tuning for CPT and IT.
Figure 5 illustrates the comparison between joint and sequential training. In both cases, different from IT-only results shown in Figure 3, we see improved performance on similar and novel tasks. This supports our hypothesis that CPT can help improve the generalization of IT, as the concepts are likely able to be shared across different tasks. It is further interesting to see that even the general tasks are improved, indicating that there could be positive transfer between CPT and IT. Comparing the two, we observe that joint training outperforms sequential training across financial and general tasks, as well as similar and novel tasks, highlighting the importance of preventing forgetting of CPT and knowledge transfer between CPT and IT.
Comparison with LoRA. In Section 3.3, we showed that LoRA can effectively adapt tasks but still suffers from task generalization. While we already showed that CPT can help in full-model training setting, we now explore whether CPT can help in the PEFT setting as well. Figure 6 presents the results of applying LoRA for IT and LoRA for both CPT and IT. Surprisingly, we find that full fine-tuning significantly outperforms the LoRA counterparts across similar and novel tasks. This finding contrasts with our previous observations in Figure 4, where performance with and without LoRA was comparable. Our results reveal that knowledge transfer from CPT to IT, which is crucial for task generalization, requires full-model training.
Takeaways from Continual Pre-training and Instruction Tuning CPT and IT have different effects and challenges. CPT is effective at introducing domain concepts. However, addressing the forgetting of general concepts and IF capability is crucial. In contrast, IT is effective at introducing domain-specific tasks, with negligible forgetting of IF capabilities and only slight forgetting in general tasks. To improve generalizability, mixing general-domain text and general tasks, along with joint training for IT and CPT is found to be more effective. Furthermore, PEFT is beneficial for task adaptation, but transferring knowledge from CPT to IT to improve task generalization requires full-model training.
3.5 Improving Reasoning with Preference Learning
In §3.2 & §3.3, we observed that with CPT and IT, the model improves in capabilities such as concepts, tasks and IF/Chat. However, such improvements were missing on the reasoning tasks. Preference Alignment (PA), where the model is trained to assign higher probability mass to better generations, has been shown to be effective in enhancing reasoning capabilities of LLMs (Pang et al., 2024; Lambert et al., 2024; Jiao et al., 2024; Wang et al., 2024). Specifically, we employ Direct Preference Optimization (DPO)  (Rafailov et al., 2023), which directly learns from positive (chosen) and negative (rejected) preference data, yielding an easier way for PA compared to reinforcement learning. We synthetically generate such data from the on-policy model, i.e., the jointly trained CPT+IT checkpoint, as it has shown the strongest performance in previous experiments.
Figure 7: Average performance on selected datasets for PA training from the ‘CPT+IT’ checkpoint on PA-Mix and PA-In.
Negligible Forgetting in PA. As with CPT and IT, we begin by performing an ablation study on different data versions to evaluate their effectiveness. Since the degree of forgetting diminishes from CPT to IT (as observed in §3.3), we expect it to be even less pronounced in PA. To quickly evaluate this hypothesis, we take a naive approach and create PA-Mix and PA-In by using either the provided or GPT4o generated responses (as done for IT in §3.3) as the ‘chosen’ samples and the output of ‘CPT+IT’ checkpoint as the ‘rejected’ ones, based on the prompts of IT-Mix and IT-In, respectively.
Figure 7 shows the results after PA training for PA-In and PA-Mix from the ‘CPT+IT’ checkpoint. [^3] PA trained from Llama3-8b-instruction has shown worse results compared to training from the ‘CPT+IT’ checkpoint in our preliminary experiments, as PA requires a strong initialization checkpoint. For instance, PA-Mix from Llama3-8b-instruction achieves only 29.99 on EDTSUM, whereas the CPT+IT’ counterpart achieves 54.21. As a results, we only investigate training PA from ‘CPT+IT’ checkpoint. We observe that PA-In performs comparably to PA-Mix, indicating that it may not be essential to include general tasks to prevent forgetting of concepts or tasks, unlike the cases of CPT and IT. This suggests that PA training can focus on in-domain tasks, without requiring a broader set of general tasks or raising concerns about forgetting. Given this, we use CFA exams (Table 4 in §3.1) as a representative source for in-domain reasoning because they cover diverse financial scenarios, emphasize complex reasoning, and, most importantly, are derived from real-world exams. These characteristics make them a strong proxy for a broader range of financial tasks, ensuring that the model generalizes effectively within the financial domain while simplifying the training process.
Another crucial observation is that there is not much difference even for unseen similar tasks (FiQA SA, FOMC and NER) and reasoning tasks (CFA-Easy and CFA-Challenge). This highlights the limitations of the current naive PA approach and suggests room for further improvement. In the following, we propose a novel PA approach that constructs preference data guided by both outcome and process reward signals.
Proposed Method and Comparison with Prior Work. Substantial efforts have been made in the past to improve reasoning with LLMs. These lead to two main families of approaches: inference-time (Snell et al., 2024; Wu et al., 2024) and training-time (Jiao et al., 2024; Qu et al., 2024) methods. Inference-time methods such as search-based (Snell et al., 2024) and self-correction (Qu et al., 2024) achieve good performance by scaling up inference-time computation. Despite their effectiveness, these methods often suffer from high latency due to the frequent assessment of intermediate reasoning states and the extensive exploration space. For example, a Monte Carlo Tree Search (MCTS) method (Jones, 2021) may improve task performance, but potentially requires significantly more computation than simply sampling solutions multiple times. This makes them unsuitable for the finance domain, where applications often require rapid responses, such as for real-time market analysis or portfolio management. In contrast, training-time methods first collect trajectories using (variations of) inference-time methods and then train the LLM with the collected trajectories. This helps the model reason more accurately and faster during inference. Therefore, in this work, we adopt the training-time approach and limit ourselves to simple CoT based inference methods.
Figure 8: An overview of the proposed Trajectory-GenORM and Trajectory-GenPRM. In Trajectory-GenORM, we collect trajectories from GenORM by evaluating the entire solution (thus question marks in the intermediate steps). In Trajectory-GenPRM, we collect trajectories from GenPRM, by identifing and correcting the first erroneous step.
To collect reasoning trajectories, there are two main approaches. The first is search-based (Setlur et al., 2024; Snell et al., 2024), where a trained Reward Model (RM) or verifier is used to guide a search method (e.g., Best-of-N, Beam Search) to identify the best reasoning path. The second is revision-based (Bai et al., 2022; Du et al., 2023; Madaan et al., 2023; Saunders et al., 2022), which attempts to improve the generation distribution through multi-round interactions, often by leveraging feedback from itself or another strong LLM to refine the input prompt.
In practice, revision-based methods have shown mixed results and have not yet been well established as reliable for achieving improvements (Huang et al., 2024a). In contrast, search-based methods have been shown to be more effective but they heavily rely on the performance of the RM. Most RMs typically assign numerical scores to candidate solutions, which do not fully exploit the text-generation capabilities that LLMs are inherently designed for. Consequently, there has been growing interest in Generative Reward Models (GenRMs), where the RM is a generative model capable of producing text-based feedback. However, most existing GenRM still either provide only outcome-level (i.e., on the final answer) signals (Mahan et al., 2024) or require extensive training (Zhang et al., 2024), which results in sparse supervision and under-utilizes the strengths of pre-trained LLMs.
To overcome these limitations, we propose leveraging the advanced text generation capabilities of LLMs as a GenRM to provide rewards on both the final answer (outcome reward) and the reasoning steps (process reward). Figure 8 illustrates the proposed method. Specifically, we construct preference data from Generative Outcome Reward Model (GenORM) and Generative Process Reward Model (GenPRM) trajectories, both of which are implemented using GPT-4o:
$\bullet$ Trajectory from GenORM (Trajectory-GenORM). Given a prompt and a candidate model generated solution, we prompt the GenORM to give a holistic judgment for the entire solution using a single ‘Yes’ or ‘No’ token. We then use the correct solutions as chosen samples and the incorrect solutions as rejected samples.
$\bullet$ Trajectory from GenPRM (Trajectory-GenPRM). Since reasoning is often complex and process rewards have been shown to be more effective (Lightman et al., 2024), we further leverage GenPRM to provide process rewards. Instead of requesting rewards at each step, which has been shown to be unnecessary by Lightman et al. (2024) and Luo et al. (2024), we prompt the GenPRM to identify the first erroneous step. While identifying the first error is helpful for training a Reward Model (RM), it is insufficient for generating preference data. To address this, we further extend the approach by prompting the GenPRM to correct the identified erroneous step.
Figure 9: Average performance on selected datasets for training from ‘CPT+IT’ on PA using Trajectory-GenORM, noted as PA (GenORM) and Trajectory-GenORM + Trajectory-GenPRM, noted as PA (GenORM+GenPRM).
Using this correction, we construct a preference data sample. The input prompt is formed by concatenating the original question, the candidate reasoning steps up to the first error, and a follow-up question framed as “What is the next step?”. The chosen response of this preference sample is the newly-obtained corrected step, while the original first erroneous step is deemed as rejected response. This approach produces trajectories that focus on predicting the correct next step given a reasoning prefix, rather than requiring a prediction of the entire reasoning trajectory, as in Trajectory-GenORM. Additional details about the prompts used for our data collection are provided in Appendix C.
Results of the Proposed Method. Figure 9 shows the results. We observe significant improvement in both CFA-Easy and CFA-Challenge, indicating that the supervision provided by Trajectory-GenORM and Trajectory-GenPRM is highly beneficial for reasoning tasks. Moreover, combining Trajectory-GenORM and Trajectory-GenPRM yields additional improvements over Trajectory-GenORM alone, demonstrating that outcome and process rewards complement each other effectively.
Takeaways from Preference Alignment PA is effective for learning to reason. Similar to IT, the forgetting of general concepts and IF capability is negligible, as is the forgetting of general tasks. However, synthesizing preference data for effective adaptation is challenging. Constructing preference data by collecting trajectories from both outcome and process signals of a GenRM is found to be an effective method.
4 Final Recipe and Evaluation
In this section, we present our final recipe based on the desired capabilities in Section 2.1, and data and model ablations performed in Section 3. A summary of the final recipe is given in Table 5. For evaluation, we compare with a wide range of baselines models, including Llama-Fin against a wide range of baseline models, including its base model, Llama3-8B-instruct, and the 8B peer, Llama3.1-8B-instruct. We also include comparisons with models of other sizes, such as Phi-3.5-mini-instruct (Abdin et al., 2024) (3.8B), and Mistral-Nemo-instruct (Jiang et al., 2023) (12B), as well as the closed model GPT-4o (OpenAI, 2023). Additionally, we evaluate against the latest open finance-specific LLM, Palmyra-Fin-32k (Writer, 2024), which is also based on the Llama3-8B-instruct model. [^4] We note that there are additional financial LLMs available, such as FinMa (Xie et al., 2023) based on Llama2, Finance-LLM (Cheng et al., 2024) based on Llama3-base and FinLLaVA (Xie et al., 2024b) focuses on multi-modality and not publicly available. However, they are either significantly smaller in scale, based on less advanced LLMs compared to our model or not publicly available. In our preliminary experiments, these models performed considerably worse than both our model and the baselines. Therefore, we have only included the SoTA financial LLM in our comparisons.

### v2

@@ Appendix C Other Capabilities
Besides those core capabilities mentioned in §3.1, domains may vary significantly in their sensitivity. For instance, the medical domain is highly sensitive, requiring utmost accuracy and strict adherence to ethical considerations. In contrast, domains such as entertainment may have more relaxed requirements. Another important consideration is multi-modality, as some domains require handling multiple types of input and output formats. For example, the healthcare domain may involve processing medical images alongside textual reports, while the e-commerce domain may integrate product descriptions, images, and customer reviews into a unified response. Similarly, scientific research often combines charts, graphs, and textual analysis to present findings effectively.
Appendix D FinTrain
Capability Domain CPT Dataset Size Reference
Concept General NaturalInstrution 100,000 Mishra et al. (2022)
PromptSource 100,000 Bach et al. (2022)
Math 29,837 Amini et al. (2019b)
Aqua 97,500 Ling et al. (2017)
CREAK 10,200 Onoe et al. (2021)
ESNLI 549,367 Camburu et al. (2018)
QASC 8,130 Khot et al. (2020)
SODA 1,190,000 Kim et al. (2022)
StrategyQA 2,290 Geva et al. (2021)
UnifiedSKG 779,000 Xie et al. (2022)
GSM8K 7,470 Cobbe et al. (2021)
ApexInstr 1,470,000 Huang et al. (2024b)
DeepmindMath 379,000 Saxton et al. (2019)
DialogueStudio 1,070,000 Zhang et al. (2023)
Finance Fineweb-Fin 4,380,000 -
Book-Fin 4,500 -
Total 10,177,294
Table D.1: Summary of curated texts. New datasets released with FinDaP are color-highlighted for emphasis.
Capability Domain Task IT Dataset Size Reference
Tasks Finance Relation Cls. FingptFinred 27,600 Sharma et al. (2022)
NER FingptNERCls 13,500 Yang et al. (2023)
FingptNER 511 Alvarado et al. (2015)
Headline Cls. FingptHeadline 82,200 Sinha et al. (2020)
Sentiment Cls. SentimentCls 47,600 Yang et al. (2023)
SentimentTra 76,800 Yang et al. (2023)
Summariz. TradeTheEvent 258,000 Zhou et al. (2021)
IF/Chat General IF/Chat SelfInstruct 82,000 Wang et al. (2022)
SlimOrca 518,000 Lian et al. (2023)
UltraChat 774,000 Ding et al. (2023)
ShareGPT 100,000 Link
Finance QA FinanceInstruct 178,000 Link
FingptConvfinqa 8,890 Chen et al. (2022)
FlareFinqa 6,250 Chen et al. (2021)
FlareFiqa 17,100 Yang et al. (2023)
Reasoning Math QA OrcaMath 200,000 Mitra et al. (2024)
MetaMathQA 395000 Yu et al. (2023)
MathInstruct 262,000 Yue et al. (2023)
Code QA MagicodeInstruct 111,000 Luo et al. (2023)
Finance CFA Exam Exercise 2,950 -
Total 3,161,401
Table D.2: Summary of our curated prompts. New datasets released with FinDaP are color-highlighted for emphasis. For datasets without formal references but only a URL, we provide their links.
Continual Pre-training Text Curation
To introduce domain concepts while preserving general concepts, we curate texts for CPT. Table D.1 summarizes the texts curation datasets. Specially, for general concepts, research has shown that a ‘small’ amount of general text (as little as 1%) can effectively mitigate the forgetting issue (Scialom et al., 2022). Therefore, we focus on collecting a relatively small but high-quality set of general-domain text. To achieve this, we use verifiable text, which is text written by humans and previously used in supervised tasks in the literature. Note that this contrasts with using unverifiable web text such as C4 (Raffel et al., 2020).
For domain concept, our goal is to collect both a large volume of data and maintain high quality. Following practices from the literature on training general LLMs (Lambert et al., 2024; Gunasekar et al., 2023), we source financial texts from primarily relevant websites and books. Specifically, we source financial text from two primary resources. The first source is web text, where we filter non-financial content from the FineWeb using URLs like ‘sec.gov’ and ‘investopedia.com’. The second source is books. We manually select 10 finance-related topics (e.g., ‘economics’ and ‘management’), download books on these topics, and convert them to text using OCR (Malmgren, 2014). Since OCR can make mistakes, we further employ a strong LLM to filter out content lacking educational value or unrelated to finance. Details on the financial URLs, finance-related topics, and the prompts used for filtering is shown below:
$\bullet$ Selected Financial URLs. We curated a selection of 70 financial websites to comprehensively cover diverse aspects of finance-related content on the web. These include trusted sources from financial institutions, regulatory agencies, educational platforms, and industry-specific news outlets. This diverse collection ensures representation across sub-domains such as investment, banking, personal finance, regulatory compliance, and financial planning, offering a well-rounded foundation that can cover most of the finance content in the web.
$\bullet$ Selected Topics. We select 12 topics that are cover most of books in finance. 5 of them are from business areas, including business, Accounting, Accounting, Management, Marketing, Trading; 1 is from Mathematics, i.e., Mathematical Economics; and 4 are from Economy area, including economy, econometrics, investing, and markets. We crawled the web to downloaded the books from the corresponding topics. For CFA, we use the material provided by CFA prep providers.
$\bullet$ Prompt for Filtering the Text. We explored various prompt formats to automatically extract an financial score using an LLM and found that the additive scale by Yuan et al. (2024) worked best. Figure D.1 shows the prompt we used to filter the ‘low-quality’ text. Specifically, this prompt allows the LLM to reason about each additional point awarded, unlike the single-rating Likert scale which fits samples into predefined boxes. Then, to avoid the LLM favoring highly technical content like academia papers, we focused on financial student level knowledge. By setting a threshold of 4 (on a scale of 0 to 5) during the filtering process, we were able to also retain some high-quality financial content.
Figure D.1: Prompt for filtering the text
Insturction Prompt Curation
Prompts represent the diverse ways users may interact with models and serves the essential component for IT and PA. Table D.2 summarizes the prompts curation datasets. Specifically, we conduct a broad survey and source general, financial, instruction-following, and reasoning tasks from public datasets. To promote diversity, we include datasets like Flare-FinQA  (Chen et al., 2021), a large open QA dataset in finance, and UltraChat (Ding et al., 2023), a dataset shown to perform well for IT in the literature (Tunstall et al., 2024; Ivison et al., 2024). Additionally, we find that exercises or demonstrations from books that were curated in §D is valuable for reasoning tasks as they usually involve challenging reasonings and come with ground truth answers and sometimes even include human-written chain-of-thought (CoT) explanations.
Figure D.2 shows the prompt we used to extract exercises from books. We carefully design the prompt to extract both the question part of an exercise, which potentially include questions, scenario and exhibits, and the answer part of the exercise, which may include answer choices and solution. In books, questions and their corresponding answers can be located far apart (e.g., the questions may appear at the beginning while the solutions are provided at the end), meaning they may not be captured within the same chunk. As a result, some questions may not have corresponding extracted answers. For such cases, GPT-4o’s generated answers are used when converting the prompt into instruction-following or preference-alignment data, as detailed in Sections B.2 and 3.2.1.
Figure D.2: Prompt for extracting exercises from books
Appendix E FinEval
@@
ESG Issue Classification MLESG 300 Chen et al. (2023a)
@@
Table E.1: Summary of our evaluation dataset. New datasets released with FinDaP are color-highlighted for emphasis.
With the breakdown of capabilities in §3.1, our evaluation framework consists of a suite for assessing these capabilities using development sets and unseen (held-out) evaluation sets. Our development set is directly split from the training data at each stage. Table E.1 outlines the capabilities and the evaluation benchmarks selected to cover these capabilities. Crucially, we did not examine scores on our unseen set while developing the models, which allows us to observe how much we may have overfitted to particular evaluations in our decisions around training recipe. For the unseen tasks (Table E.1), we manually review each individual dataset and have the following considerations.
@@
$\bullet$ Benchmarking tasks. Corresponding to the capabilities, we consider a diverse set of benchmarking tasks. For concepts, we include knowledge tasks in the general domain, such as AI2-ARC (Clark et al., 2018), as well as in finance, such as MMLU-Finance (Hendrycks et al., 2021). For tasks, we consider general tasks, such as Social-IQA (Welbl et al., 2017), and domain-specific tasks, such as MLESG (Chen et al., 2023a). Notably, we intentionally include a few financial tasks such as Flare-TATQA (Zhu et al., 2021) and SM-Bigdata (Soun et al., 2022) that require understanding of tabular data, as this data format is common in this domain. For IF/Chat capabilities, we utilize popular instruction-following benchmarks, such as MT-Bench (Zheng et al., 2023). For reasoning, we include general reasoning tasks, such as MathQA (Amini et al., 2019a) and Hellaswag common sense reasoning (Zellers et al., 2019), as well as domain-specific reasoning tasks, such as CRA-ProroSeguro claim analysis (Feng et al., 2024). We also construct a new benchmark on CFA-Challenge based on CFA Level III, one of the most challenging financial exams that requires comprehensive reasoning (Khamnuansin et al., 2024; Callanan et al., 2024).
@@
$\bullet$ Evaluation method. We split our evaluation set into two types based on their exposure to Instruction tuning (IT) data (Table E.1). The first type, Similar, includes tasks whose types have been encountered during training, even if the specific tasks themselves are unseen (e.g., a new NER task). The second type, Novel, includes tasks whose types have not been seen during training, representing entirely new challenges for the model (e.g., stock movement prediction). We use two different evaluation methods based on the nature of the benchmarks. For knowledge and NLP tasks (e.g., NER), we employ a straightforward direct answer evaluation. For reasoning tasks (e.g., CFA-Challenge), we use a 0-shot chain-of-thought (CoT) (Wei et al., 2023) answer evaluation to enhance the reliability of our evaluation. This also exposes the reasoning path, allowing us to investigate the causes of incorrect answers and enabling a more fine-grained comparison across different models.
Appendix F Summary of the Final Recipe and Hyper-parameters